How do you take a negative fraction to a negative power? Ex.) -2/3 to the -3 power
Answer by lwsshak3(6463)

How do you take a negative fraction to a negative power? Ex.) -2/3 to the -3 power
It depends on how you write it.
If you write it this way:
(-2/3)^-3
=1/(-2/3)^3
=1/-8/27
=-27/8
..
If you write it this way:
-(2/3)^-3
=-1/(-2/3)^3
=-1/-8/27
=27/8